solve
-7X - 23= -25 1/3
x=what

Answer by ewatrrr(24785)


Hi,
To solve this equation, we must isolate x on one side of the equation.
This is done by performing the same operations to both sides of the equation,
-7x - 23 = -25 1/3 Adding 7x to both sides
-23 = 7x -25 1/3 Adding 25 1/3 to both sides
-23 + 25 1/3 = 7x
2 1/3 = 7x simplifying 2 1/3 = 7/3
7/3 = 7x dividing both sides by 7
1/3 = x